Grade 8Mechanics

a string of negligible thickness is wrapped several times around a cylinder kept on a rough horizontal surface.a man standing at a distance l from the cylinder holds on end of the string and pulls the string and pulls the cylinder towards him.there is no slipping anywhere.the length of the string passed through the hand of man while the cylinder reaches his handis??

mgayathri
9 Years agoGrade 8

1 Answer

Abhishek
8 Years ago
Suppose radius is rWhen cylinder moves a distance L it swepts angle L/rSo extra string unwound =L/r×r=LSo total string to be pulled is L+L=2L